What is the inverse of this equation F(X)=3x/5+3

Answer by stanbon(75887)

What is the inverse of this equation F(X)=3x/5+3
--------
1st; Interchange x and y:
x = 3y/5+3
-----
2nd: Solve for "y":
3y/5 = x-3
------
4y = 5(x-3)
------
y = (5/4)(x-3)
-----
That is the inverse.
Cheers,
Stan H.
====================